Klapperich Lab:Notebook/Lab Meeting Notes/2010/03/18

From OpenWetWare
Jump to navigationJump to search
Owwnotebook icon.png Project name Report.pngMain project page
Resultset previous.pngPrevious entry      Next entryResultset next.png

18 March 2010 Lab Meeting

Jason Keller Presentation.

‡ Announcements

‡ Flu R01:Integration

  • meetings every other M 1-3pm. 705.

† HDA (Lead: Jaephil, Team:Sonali)

  • Patient Samples - dilution experiments (MM/SH)
  • Start planning R01 for Submission on 6/5/10. MM, CMK.
  • HDA mix in dried form - Expecting it from BH -- ON HOLD.
  • warm start on the way - MTA (CMK)
  • 2nd HDA design - "microSPE" + "Flu primer" + "HDA in a cup" + "Detection kit from Biohelix"

† Sample Concentration (Lead: Jane, Team: Jaephil)

  • Y channels working better. Try virus in PDMS, Make COP version of chip.
  • Samples expected from Fauchet group. Contact today.

† SPE Column Optimization for DNA/RNA. (Lead: Sonali, Team: Hussam, Jessie, Brendan)

  • (In progress) Jessie will run chips with serial virus dilutions from 1:1 to 1:1,000,000 to re-do previous dilution experimnent with WHO SYBR green assay on chip.
  • Less input sample (same concentration, 37.5uL sample vs 100uL sample). From raw data (from 30uL extracted RNA), less sample gives significantly(p-value = 0.0005) less copies. If normalized to per mL of sample input, copies/mL are comparable between two sample inputs (p-value = 0.5094).(JC)
  • Plaque assay on patient samples for QQ's paper.
  • Casein in human serum experiments ongoing. (HM)

† Integrated chip for flu(SPE+RT+PCR) (Lead: Qingqing)

  • Real copy number of previous human samples. So QQ is working with the higher concentration samples. Range is 10^6-10^8 of copy number/ml.
  • REVISIT the % recovery number.
  • QQ will try: more enzyme, Tween 20, BSA amounts.

  • PCR1 will combine with PCR2 if PCR2 comes back for revisions. (CMK)

‡C. diff Project (Cathie, Sonali, Satish Singh, Lisa J., His post doc )

  • See email notes.

‡ Coulter Flu Fraunhofer Project (Lead: Sonali, Team: Sonali, Jessie, Cathie, CMI Folks, Qingqing)

  • New meeting time, T, 9:30 am every other week. MEETINGS CANCELLED THIS WEEK.

‡ Agilent Automated Sample Preparation (Lead: Alex)

  • Paper on HotDog - with CMK/ASB

‡ COBRA (Lead: Jaephil, Team: Cathie, Jane for virus only, PHO folks)

  • Evap paper - Additional experiment for Fig 3. By 3/15.

‡ Biointerfaces group (Lead: MinCheol, Team: Cathie, MCK, Wong and Meller folks)
* Cathie will submit paper.

‡ CIMIT- Sepsis (Lead:Cathie, Team:TBA)

  • Sample collection ongoing.

‡ PATH Grant (Lead:Cathie, Team:Frank, Sean, Mark, Jake Trueb (ME))

  • Phone call with PATH 3/22. Frank has class 2-4 but will come ASAP.
  • Frank: One week data up 3/5/10. Starting Friday will do saturation/capacity experiment. Longer-term time points with current protocol will be run this weekend.
  • Sean: Started making straws for the new experiments we talked about (DNA weights of 1ng, 10ng, and 100ng through 700nm at SPE 100 monolith at 0-day time point).
  • Mark: as of 3/17/10 up to 10% recovery from 10 ng input sample with water elution. Also up to 13% recovery with 100ng/straw. Loaded two sample groups for multiple temperatures for 1-week and 2-week/1-month time points. 1-week trial completed as of 3/24. Results peak at 20% (-70C) and decrease with temperature. Next to test straws with Qiagen buffer.

‡ IIH Senior Project.

  • Progress Report 2 due today
  • Finished image and data analysis of 100 micrometer films.
  • Working on 44 micron films.
  • Drafting ELISA.